OpenStudy (anonymous):

x-1/x^2-1 what is the hole, x and y intercept?

OpenStudy (anonymous):

factor and cancel

OpenStudy (anonymous):

would the hole be (1,0)?

OpenStudy (anonymous):

thanks!

OpenStudy (yb1996):

to find the y-intercept, all you have to to is plug in 0 for x

OpenStudy (anonymous):

i don't think it is \((1,0)\)

OpenStudy (anonymous):

oh is it (1,1/2)?

OpenStudy (anonymous):

\[\frac{x-1}{(x+1)(x-1)}=\frac{1}{x+1}\] if you put \(x=1\) in \(\frac{1}{x+1}\) you get \(\frac{1}{2}\)

OpenStudy (anonymous):

Is there any x-intercept I keep getting zero?